OpenStudy (anonymous):

Write the equation in slope-intercept form of the line that has a slope of 2 and contains the point (3, 7).

OpenStudy (blurbendy):

slope-intercept form: y = mx + b Use: y - y1 = 2 (x - x1) substitute (3, 7) in for x1 and y1 in the above equation to get it in the form y = mx + b understand?